OpenStudy (ray1998):

Please help? Fan and medal will be given. When radioactive substances decay, the amount remaining will form a geometric sequence when measured over constant intervals of time. The table shows the amount of a radioactive isotope initially and after 2 hours. What are the amounts left after 1 hour, 3 hours, and 4 hours? Hours Elapsed - 0, 1, 2, 3, 4 Grams of the isotope - 1986, ___, 477, ___, ___ How much of the isotope is left after 1 hour? How many grams? Thank you so much in advance!

OpenStudy (palwandew):

Is -1986 is amount at initial time?

OpenStudy (ray1998):

Sorry, not negative, it's positive. I just put a slash there to indicate how many hours were elapsed.

OpenStudy (palwandew):

Here a1= 1986, a3=477 and r=?. Right

OpenStudy (ray1998):

I think so. Yes :) and I apologize for the late reply.

OpenStudy (palwandew):

To find "r" use the formula an=a1*r^n-1

OpenStudy (ray1998):

Okay. So, 1986*r^n-1

OpenStudy (palwandew):

Here we have a1=1986 and a3=477 we start with a3=a1*r^3-1 477=(1986)*r^2 477/1986=r^2 0.24=r^2 \[\sqrt{0.24}\]=r r=0.48

OpenStudy (palwandew):

Now to find a2 Use a2=a1*r^2-1

OpenStudy (ray1998):

1986 *0.48^2-1 = 953.28?

OpenStudy (ray1998):

Here are the terms: 1986, 953, 477, 219, 105. How do I find out how much of the isotope is left after 1 hour and how many grams are left?

OpenStudy (anonymous):

you got the 953 somehow right? not sure how you got it, but you got it

OpenStudy (ray1998):

Yes :) I got it by multiplying a1 by the common ratio. 1986 * 0.48

OpenStudy (anonymous):

that is not what i get

OpenStudy (anonymous):

let me try again, maybe i made a mistake

OpenStudy (anonymous):

\[1986r^2=477\\ r^2=\frac{477}{1986}\\ r=\sqrt{\frac{477}{1986}}\]

OpenStudy (ray1998):

953.28 is the actual answer I get when multiplying the two.

OpenStudy (anonymous):

which is close to \(0.48\) but i get \(0.49\)

OpenStudy (anonymous):

anyway you have answered the question assuming these numbers are right (i didn't check them) you have \[ 1986, 953, 477, 219, 105. \]

OpenStudy (ray1998):

Yes

OpenStudy (anonymous):

first is at 0 hour, second is one hour, third is two hour, etc

OpenStudy (ray1998):

Okay